OpenStudy (anonymous):

-3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this

OpenStudy (anonymous):

Name your equations -3x + 7y = -16 i) -9x + 5y = 16 ii) i) x3

OpenStudy (anonymous):

what do u get?

OpenStudy (anonymous):

-9x+21=-48 iii) iii)- ii)

OpenStudy (anonymous):

sorry meant to say -9x+21y=-48 iii) iii)- ii)

OpenStudy (anonymous):

16y=-64 y=-4

OpenStudy (anonymous):

sub y into ii) -9x + 5(-4) = 16 -9x=36 x=-4

OpenStudy (anonymous):

check the answer by sub x and y values into any of the equations